OpenStudy (anonymous):

The length of a rectangular boat storage shed must be 4 meters more than the width, and the perimeter must be at least 116 meters. What is the range of values for the width?

OpenStudy (anonymous):

Possible Answers: A. Exactly 27 meters B. At most 27 meters C. At least 27 meters

OpenStudy (anonymous):

L = W +4 Perimeter = 2L + 2W 116=2L+2W 116=2(W+4)+2W now you need to solve for W

OpenStudy (anonymous):

Ok. So doing 116=2(W+4)+2W the answer would be exactly 27 meters.